OpenStudy (iwillrektyou):

Can someone please help me on a brain teaser/algebra 1 problem? Will medal best helper! It's supposed to be quite simple, but I kinda hit my head earlier so I'm kinda groggy XD (Question will be pasted in next message.)

OpenStudy (iwillrektyou):

The problem is: Three boys, Josh, Hans, and Simon, are sitting around on a farm with nothing to do. They walk into the barn and notice a scale used to weigh cattle. They decide to weigh themselves. Unfortunately, the scale begins at 100 kg. The problem is that none of the boys weigh more than 100 kg. They decide to weight themselves in pairs. Josh was sure that he weighed the most. Hans and Simon together = 132 kg Simon and Josh together = 151 kg Josh and Hans together = 137 kg What do they each weigh?

OpenStudy (iwillrektyou):

I know you have to make an algebraic equation, and combine like terms, but I don't know how to start.

jimthompson5910 (jim_thompson5910):

H = Han's weight S = Simon's weight J = Josh's weight `Hans and Simon together = 132 kg` H+S = 132 `Simon and Josh together = 151 kg` S+J = 151 `Josh and Hans together = 137 kg` J + H = 137

jimthompson5910 (jim_thompson5910):

So the three equations in this system are H+S = 132 S+J = 151 J + H = 137

OpenStudy (iwillrektyou):

Ok.

jimthompson5910 (jim_thompson5910):

Solve equation 1 for S to get H+S = 132 H+S-H = 132-H S = 132-H Then plug this into equation 2 S+J = 151 (132-H)+J = 151 132-H+J = 151

jimthompson5910 (jim_thompson5910):

So we now have 132-H+J = 151 J + H = 137

jimthompson5910 (jim_thompson5910):

hopefully this is making sense?

OpenStudy (iwillrektyou):

It is. So basically you isolate each variable by replacing the other and solving for it?

jimthompson5910 (jim_thompson5910):

yeah a bunch of substitutions really

OpenStudy (iwillrektyou):

Ok thank you.

jimthompson5910 (jim_thompson5910):

no problem